### v1

@@ 4.7 Pathology Report Generation
The task of pathology report generation is inspired by existing works on Chest X-ray and other medical report generation [104, 105, 106]. In this task, the report generation model takes a WSI as input and generates the corresponding pathology report. Specifically, the input WSI is first processed by foundation models to extract an initial representation. This representation is then fed into the encoder-decoder architecture of report generation models to produce the decoded pathology report. During this process, the visual encoder further processes the initial representations of WSIs through specific designs [105, 106, 63] to obtain the optimal WSI features for the report decoding stage. The text decoder of the model then utilizes these features for report generation. A good initial representation of WSI could significantly facilitate both the visual encoding and textual decoding stages. Consequently, the quality of the generated report is directly influenced by the representations provided by the foundation models. In this task, we adopt the HistGen model [63] for WSI report generation and set the learning rate to 1e-4, weight decay to 0.8 per epoch. The model is trained for 40 epochs with batch size 1 using features extracted from different foundation models.
@@ 6 Code availability
The code and weights of the GPFM will be made available upon acceptance.

### v2

@@ 4.7 Pathology Report Generation
The task of pathology report generation is inspired by existing works on Chest X-ray and other medical report generation [104, 105, 106]. In this task, the report generation model takes a WSI as input and generates the corresponding pathology report. Specifically, the input WSI is first processed by foundation models to extract an initial representation. This representation is then fed into the encoder-decoder architecture of report generation models to produce the decoded pathology report. During this process, the visual encoder further processes the initial representations of WSIs through specific designs [105, 106, 63] to obtain the optimal WSI features for the report decoding stage. The text decoder of the model then utilizes these features for report generation. A good initial representation of WSI could significantly facilitate both the visual encoding and textual decoding stages. Consequently, the quality of the generated report is directly influenced by the representations provided by the foundation models. In this task, we adopt the HistGen model [63] for WSI report generation and set the learning rate to 1e-4, and weight decay to 0.8 per epoch. The model is trained for 40 epochs with batch size 1 using features extracted from different foundation models.
@@ 6 Code availability
The code and weights of the GPFM will be made available upon acceptance. The code and weights of the GPFM have been made available on GitHub (https://github.com/birkhoffkiki/GPFM/).
7 Ethics declarations
This project has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC). The protocol number is HREP-2024-0212.